# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

### v1

@@ 4. Methodology
Our approach consists of two phases: a systematic literature review, and a survey study. Additionally, we also conducted a complimentary analysis of broader trust literature to provide further context and insights.
@@ 4.1. Phase1: Systematic Literature Review
We conducted an SLR following an approach by Kitchenman et al.  (Kitchenham, 2012). Before starting our review process, we developed RQs according to Kitchenman’s guidelines. This helped us to systematically find papers that are related to our research goals. We then performed the following steps: 1) Search for Primary Studies 2) Screening 2a) Filter by title 2b) Filter by abstract 2c) Full-text Screening 3) Snowballing 4) Manual Addition 5) Data Extraction and Analysis. Fig.1 shows a visual representation of our complete pipeline. Each step in this process was conducted independently by the two authors, who met after each step to ensure alignment and resolve any differences. The third author was consulted as needed to help settle any disputes.
[…]
[PRISMA flow diagram]A PRISMA  (Page et al., 2021) flow diagram describing the stages of our review with the number of records at each stage
Figure 1. A PRISMA  (Page et al., 2021) flow diagram that describes our review process, labeled with the number of records at each stage
@@ 4.1.1. Search for Primary Studies
We focused on the period from January 1st 2002 to November 1st 2024. We chose 2002 as our starting period as this was when the method to model language using feed-forward neural networks was published by Bengio et al.  (Bengio et al., 2000).
Databases: We selected the following databases, including ACM Digital Library, IEEE Xplore, Springer Link, Google Scholar, and DBLP. We then formulated a search string to query these databases. These databases were chosen as they were the major venues in SE, Machine Learning, HCI, and DL that encompass our research goals.
@@
Search strategies: We used different mechanisms to query each database. For IEEEXplore and Springer Link, we used their advanced search feature to filter by date and venues and used their website to export the first 1,000 most relevant results as an Excel sheet. However, DBLP, Google Scholar, and ACM Digital Library do not have an export feature; thus, we wrote a simple Python program to query using a REST API and appropriate filters. After searching the databases with specified search strings and eliminating duplicates, we had 4,472 potentially relevant papers from this step. The script and the breakdown of the number of papers retrieved from each database are available in our repository  (AnonymousRepoTrustRepo, 2024).
@@ 4.1.2. Filtering by inclusion criteria
We had two iterative inclusion criteria to filter out papers from the primary search.
  • Filter by title: We screened all search outcomes by reviewing the title first. In this step, any papers with titles that contained keywords ”trust” OR ”distrust” OR ”trustworthiness” OR ”SE” OR ”Software Engineering” OR ”Software Development” OR ”Code Generation” OR ”Program Repair” OR ”Automatic Program Repair” OR ”Software Testing” OR ”Test Generation” OR ”LLM” OR ”Large Language Model” OR ”LLMs” OR ”Large Language Models” OR ”Deep Learning” OR ”Machine Learning” were considered for further screening. We kept all these keywords to ensure we did not lose any relevant papers. We were left with $3,186$ papers for further screening from this step.
  • Filter by abstract: Filtering by abstract was a manual effort by two authors of this paper. The abstract was systematically analyzed to confirm their relevance to our research goals. Papers that contained some indication of using trust or trust-related concepts were included in this step. Specifically, we searched for specific indicators in the abstracts, such as explicit mentions of trust, trustworthiness, or related concepts such as reliability, confidence, or user acceptance in the context of LLMs applied to SE tasks. Papers that merely mentioned LLMs or SE without a clear focus on trust were excluded. Additionally, abstracts that discussed trust in general AI or machine learning contexts without specific application to LLM4SE were also filtered out. From this step, we filtered out 3,163 papers, leaving us with 23 papers. The high number of excluded papers reflects the novelty of trust research specifically in the LLM4SE domain. Many papers may have touched on related topics but lacked the specific intersection of trust, LLMs, and software engineering that our study aimed to explore.
@@ 4.1.3. Full-text Screening
We conducted a thorough full-text review of the $23$ papers that passed the abstract filtering process. This step involved a detailed examination of each paper’s content to ensure its relevance to our research questions. We specifically looked for papers that addressed the intersection of trust-related concepts, SE, and LLMs. Papers were included if they:
@@
  • Addressed challenges or proposed solutions for building trust in LLM-assisted SE processes.
Papers that only tangentially mentioned trust or LLMs without substantial focus on SE applications were excluded at this stage.
@@ 4.1.4. Snowballing and Manual Addition of Studies
To account for studies we might have missed with our primary search keywords, we performed snowballing on the resulting studies from our inclusion criteria. Papers in the references went through the same inclusion criteria as papers from our primary study. We first filtered them by title, and then by abstract. We added 6 papers from this step.
@@ 4.1.5. Exclusion criteria
In this step, we applied exclusion criteria to filter out studies that were irrelevant to our research questions. This involved a manual review of all papers to ensure their applicability. While trust has been widely studied in DL and HCI, we excluded papers that did not explicitly operationalize trust in SE. Our focus is on understanding trust in the context of SE, particularly regarding the integration and use of LLMs. Although trust research in DL and HCI offers valuable insights, it does not directly address SE-specific challenges such as workflow integration, developer-user interactions, and task-specific trust calibration. Therefore, we exclude these papers from our main analysis. However, to enrich our analysis, we incorporate relevant findings from broader trust research that, while not explicitly focused on SE, provide conceptual insights applicable to our study (see 4.3 for the methodology for selecting these articles). The finalized corpus contains 18 papers.
@@ 4.1.6. Data Extraction and Analysis
We iteratively went through collected papers to collect both general and specific information relevant to our RQs, following protocols suggested by  (Levett, 2022) to extract data. We create a dataset with the following data: authors, year of publication, RQs proposed, methodology, study population and characteristics (for user studies), the definition of trust, consequences of trust, the importance of trust, antecedents of trust, and metric used for trust.
Both authors worked independently to extract information and fill out the database. The authors then discussed any discrepancies with the third author. In the end, we obtained a high Inter-Rater Reliability (IRR), with a Cohen’s kappa(k) coefficient of 0.824.
Based on the extracted information, we derive our taxonomy to answer our RQs. We draw upon the framework and vocabulary from the MATCH model (Liao and Sundar, 2022) to derive our taxonomy to answer RQ2 and RQ3. The MATCH model presents factors that influence a user’s trust throughout the AI interaction process, deriving these insights from a comprehensive review of the literature, making it appropriate for our study. The MATCH model breaks down the trust-formation process into three main components:
@@
    Model Trustworthiness Attribute: These attributes are derived from the ABI framework (Dietz and Hartog, 2006) which includes three critical qualities. Ability refers to what the model can do, showcasing its competence and performance in SE-related tasks. Intention Benevolence addresses why the model was made, reflecting the intentions behind its development. Process Integrity involves how the model was trained, ensuring adherence to ethical standards and that the training data is free of biases.
@@
    Trust Affordance: Trust affordance refers to the exhibited characteristics of a system that indicate how it can be used, guiding users in evaluating the model’s trustworthiness. Trust affordance can manifest in three ways. AI-generated content is prediction, answers, and other outputs that can directly indicate the AI’s ability. Transparency regarding the model’s inner functioning can enhance users’ understanding of the model’s benevolence and process integrity. Interaction covers the ways users interact with AI, which can influence trust through the usability and design of the system.
@@
    Trust Processing: Based on users’ ability and motivation, they will partake in two different procedures of trust judgment-making. Systematic Processing involves a rigorous assessment of information to form rational trust judgments. Heuristic Processing: uses cognitive shortcuts to make quicker trust judgments, sometimes leading to incorrect judgment.
@@ 4.2. Phase2: Survey Study
With the approval of our Institutional Review Board (IRB), we conducted a survey study to investigate practitioners’ perceptions of trust-related concepts when using LLM4SE tasks. We chose a survey study over interviews or mixed methods to gather a broader and more diverse range of practitioner insights on trust in LLMs for SE tasks. User studies, particularly surveys, enable efficient data collection from larger samples while maintaining consistency for comparisons between participant groups. Unlike interviews, which are time-intensive and prone to interviewer bias, surveys provide anonymity, reducing social desirability bias, and ensuring honest responses. Additionally, the inclusion of open-ended questions in the survey allowed us to capture qualitative insights alongside quantitative data, offering a balance between depth and scalability suitable for our exploratory research goals. This approach aligns with established practices in trust research, improving the relevance and generalizability of our findings. In this study, we specifically explore the definitions of trust, its antecedents, and metrics, complementing our systematic review of the literature with insights from both experienced and novice practitioners. We employed purposive sampling  (Baltes and Ralph, 2021) to reach out to practitioners with varying levels of SE and machine learning expertise, including industrial experts, researchers, students, and software engineers. Participation was voluntary, with no incentives offered, and participants were informed about the voluntary nature of our study during solicitation.
@@ 4.2.1. Survey Structure
Section 1 examines the practitioner’s familiarity with using LLMs for various SE tasks. It qualifies practitioners for the rest of the survey, ensuring valid feedback by excluding those who have never used LLMs for the subsequent downstream tasks.
Section 2 combines blocks of questions investigating practitioners’ insights into trust-related concepts in Test Case Generation, Code Generation, and Program Repair. Blocks are presented in random order to avoid bias. Each block covers trust definitions, antecedents, and metrics. Only practitioners with experience in using LLMs for these tasks answer these questions.
Section 3 captures the general perception of the importance of trust, the desired metric for trust, and whether practitioners value the importance of these topics.
@@ 4.2.2. Qualitative Metric
  • Definition: We ask practitioners with experience in using LLMs for specific tasks to define trust in code generation, test-case generation, and program repair.
  • Antecedents of Trust: We gather practitioners’ views on useful trust antecedents, as explained in  4.2.3, for particular downstream SE tasks, which allows us to investigate if trust antecedents are consistent across downstream tasks.
  • Trust Metrics: We collect information on how practitioners measure trust in specific downstream tasks through open-ended questions. This helps us compare practitioners’ trust measurement methods with those in the literature. We also ask about their code review process for both human-written and LLMs-generated code to identify similarities and differences.
@@ 4.2.3. Trust Antecedents Selection
The survey asks participants to select trust antecedents for each downstream task. We identified model-related and community-related attributes from SE, HCI, and ML literature  (Li et al., 2023b; EUw, [n. d.]; körber_2018; Widder et al., 2021a; Johnson et al., 2023; Ji et al., 2024). One author compiled all the documented antecedents, and then two authors independently categorized them, resolving disagreements through discussion. We selected the nine most popular antecedents for our survey, allowing participants to add any additional ones, which were not originally listed.
@@ 4.2.4. Data Collection
We reached out to $96$ potential participants with varying expertise and from different backgrounds, who were not involved in or aware of the purpose of this work. Out of these groups, $48$ participants answered the survey. However, we had to discard the $23$ participants due to the incomplete or poor quality of their responses leaving $25$ valid responses. The study was performed using Qualtrics  (noa, [n. d.]).
@@ 4.2.5. Survey Participants
The survey included 25 participants representing a diverse range of professional roles in software engineering and research. The majority identified as students (40%), followed by researchers (20%), engineers (16%), and professors (12%), with the remaining roles distributed among freelancers, managers, and other titles. Participants demonstrated varied levels of experience in software engineering: 44% reported more than 5 years of experience, 24% had 3–5 years, and 28% had 1–2 years, with only one respondent (4%) having less than a year of experience. In terms of education, the majority held graduate or professional degrees (60%), while 36% had bachelor’s degrees, and 4% had completed high school or equivalent.
When asked about their experience using LLMs for software engineering tasks, 36% of respondents reported 1–2 years of experience, another 36% had less than 1 year, and 16% had 3–5 years of experience. This distribution reflects a relatively early adoption stage of LLMs within software engineering, aligning with the novelty of the technology in this domain. Please refer to our online appendix  (AnonymousRepoTrustRepo, 2024) for more details.
@@ 4.2.6. Survey Validity
To solidify our survey, we conducted two pilot studies: one with $5$ participants and the other with $2$ participants who were not invited to the main survey. Based on the results of these pilot studies, we eliminated any leading questions and minor errors and enhanced our selection of trust antecedents. Our survey study is also driven by the findings from our literature survey. In addition, two of the authors carried out the data analysis individually to avoid bias and human error.
@@ 4.3. Complementary Analysis of Broader Trust literature
Our initial review of trust in LLMs in SE yielded only $18$ papers, revealing significant gaps in existing literature in LLM4SE. Thus, to supplement our systematic review and provide additional context for trust in LLM4SE, we conducted a complementary analysis of broader trust literature from HCI, DL, and automation. These fields provide well-established definitions, influencing factors, and evaluation metrics that can inform our understanding of trust in LLM4SE. Thus, we extracted trust-related concepts such as definitions, influencing factors, and evaluation metrics that could inform our findings within the SE domain.

### v2

@@ 4. Methodology
Our approach consists of two phases: a systematic review of the literature and a survey study. To provide more context and information, we also conducted a complementary analysis of the broader trust literature.
@@ 4.1. Phase1: Systematic Literature Review
We conducted a Systematic Literature Review (SLR) following the approach of Kitchenman et al. (Kitchenham, 2012). Before starting our review process, we developed RQs according to Kitchenman’s guidelines. This helped us to systematically find papers that are related to our research goals. We then performed the following steps: 1) Search for Primary Studies 2) Screening 2a) Filter by title 2b) Filter by abstract 2c) Full-text Screening 3) Snowballing 4) Manual addition 5) Data Extraction and Analysis. Fig. 2 shows a visual representation of our complete pipeline. Each step in this process was conducted independently by the two authors, who met after each step to ensure alignment and resolve any differences. The third author was consulted as needed to help settle any disputes.
Figure 2. A PRISMA (Page et al., 2021) flow diagram that describes our review process, labeled with the number of records at each stage: PRISMA flow diagram
@@ 4.1.1. Search for Primary Studies
We focus on the period from January 1, 2002, to November 1, 2024. We chose 2002 as our starting period because this was when the method for modeling language using feedforward neural networks was published by Bengio et al. (Bengio et al., 2000).
Databases: We selected the following databases, including ACM Digital Library, IEEE Xplore, Springer Link, Google Scholar, and DBLP. We then formulate a search string to query these databases. These databases were chosen because they were the main venues in SE, Machine Learning (ML), HCI, and DL that encompass our research goals.
@@
Search strategies: We used different mechanisms to query each database. For IEEEXplore and Springer Link, we used their advanced search feature to filter by date and venue and used their website to export the first 1,000 most relevant results as an Excel sheet. However, DBLP, Google Scholar, and ACM Digital Library do not have an export feature; thus, we wrote a simple Python program to query using a REST API and appropriate filters. After searching the databases using the specified search strings, we retrieved 2,000 papers from ACM, 1,000 from IEEE, 460 from Google Scholar, 1,000 from Springer, and 12 from DBLP, resulting in a total of 4,472 papers. The comparatively low number of results from DBLP is expected due to its search functionality, which performs a literal match on all search terms and does not support the phrase or proximity searches available in the other databases we queried. After eliminating 232 duplicates, we had 4,240 potentially relevant papers from this step. The script and the breakdown of the number of articles retrieved from each database are available in our repository (AnonymousRepoTrustRepo, 2024).
@@ 4.1.2. Filtering by inclusion criteria
We had two iterative inclusion criteria to filter out articles from the primary search.
  • Filter by title: We selected all search results by reviewing the title first. In this step, any papers with titles that contained keywords ”trust” OR ”distrust” OR ”trustworthiness” OR ”SE” OR ”Software Engineering” OR ”Software Development” OR ”Code Generation” OR ”Program Repair” OR ”Automatic Program Repair” OR ”Software Testing” OR ”Test Generation” OR ”LLM” OR ”Large Language Model” OR ”LLMs” OR ”Large Language Models” OR ”Deep Learning” OR ”Machine Learning” were considered for further screening. We kept all these keywords to ensure that we did not lose any relevant papers. We were left with $3,186$ papers for further screening from this step.
  • Filter by abstract: Filtering by abstract was a manual effort of the two authors of this paper. The abstract was systematically analyzed to confirm its relevance to our research goals. In this step, papers that contained some indication of using trust or concepts related to trust were included. Specifically, we searched for specific indicators in the abstracts, such as explicit mentions of trust, trustworthiness, or related concepts such as reliability, confidence, or user acceptance in the context of LLMs applied to SE tasks. Papers that merely mentioned LLMs or SE without a clear focus on trust were excluded. Furthermore, abstracts that discussed trust in general AI or machine learning contexts without specific application to LLM4SE were also filtered out. From this step, we filtered out 3,163 papers, leaving us with 23 papers. The high number of excluded articles reflects the novelty of trust research, specifically in the LLM4SE domain. Many articles may have touched on related topics, but did not have the specific intersection of trust, LLM, and software engineering that our study aimed to explore.
@@ 4.1.3. Full-text Screening
We conducted a thorough full-text review of $23$ articles that passed the abstract filtering process. This step involved a detailed examination of each paper’s content to ensure its relevance to our research questions. We specifically looked for papers that addressed the intersection of trust-related concepts, SE, and LLMs. Articles were included if they:
@@
  • Addressed challenges or proposed solutions to build trust in LLM-assisted SE processes.
Papers that only tangentially mentioned trust or LLMs without a substantial focus on SE applications were excluded at this stage.
@@ 4.1.4. Snowballing and Manual Addition of Studies
To ensure a comprehensive search, we performed backward snowballing by systematically screening the reference lists of all studies that met our inclusion criteria. The articles in the references went through the same inclusion criteria as the articles in our primary study. We first filtered them by title and then by abstract. We added $6$ papers from this step.
@@ 4.1.5. Exclusion criteria
In this step, we applied exclusion criteria to filter out studies that were irrelevant to our research questions. This involved a manual review of all papers to ensure their applicability. Although trust has been widely studied in DL and HCI, we excluded papers that did not explicitly operationalize trust in SE. Our focus is on understanding trust in the context of SE, particularly regarding the integration and use of LLMs. Although trust research in DL and HCI offers valuable insights, it does not directly address SE-specific challenges such as workflow integration, developer-user interactions, and task-specific trust calibration. Therefore, we exclude these papers from our main analysis. However, to enrich our analysis, we incorporate relevant findings from broader trust research that, while not explicitly focused on SE, provide conceptual insights applicable to our study (see 4.3 for the methodology for selecting these articles). The finalized corpus contains 18 papers.
@@ 4.1.6. Data Extraction and Analysis
We iteratively reviewed the articles collected to collect general and specific information relevant to our RQ, following the protocols suggested by (Levett, 2022) to extract data. We create a dataset with the following data: authors, year of publication, proposed RQs, methodology, study population and characteristics (for user studies), definition of trust, consequences of trust, importance of trust, antecedents of trust, and metric used for trust.
Both authors independently worked to extract information and fill out the database. The authors discussed any discrepancies with the third author. In the end, we obtained a high Inter-Rater Reliability (IRR), with a Cohen’s kappa(k) coefficient of 0.824.
Based on the extracted information, we derive our taxonomy to answer our RQs. We draw on the framework and vocabulary of the MATCH model (Liao and Sundar, 2022) to derive our taxonomy to answer RQ2 and RQ3. The MATCH model presents factors that influence a user’s trust throughout the AI interaction process, deriving these insights from a comprehensive review of the literature, making it appropriate for our study. The MATCH model breaks down the trust-formation process into three main components:
@@
    Model Trustworthiness Attribute: These attributes are derived from the ABI framework (Dietz and Hartog, 2006) which includes three critical qualities. Ability refers to what the model can do, showcasing its competence and performance in SE-related tasks. Intention Benevolence addresses why the model was made, reflecting the intentions behind its development. Process Integrity involves how the model was trained, ensuring adherence to ethical standards, and that the training data are free from biases.
@@
    Trust Affordance: Trust affordance refers to the exhibited characteristics of a system that indicate how it can be used, guiding users in evaluating the model’s trustworthiness. Trust affordance can manifest itself in three ways. AI-generated content is prediction, answers, and other outputs that can directly indicate the AI’s ability. Transparency regarding the model’s inner functioning can enhance users’ understanding of the benevolence of the model and the integrity of the process. Interaction covers the ways users interact with AI, which can influence trust through the usability and design of the system.
@@
    Trust Processing: Based on users’ ability and motivation, they will partake in two different trust judgement making procedures. Systematic Processing involves a rigorous assessment of information to form rational trust judgments. Heuristic Processing: uses cognitive shortcuts to make faster trust judgments, sometimes leading to incorrect judgment.
@@ 4.2. Phase2: Survey study
With the approval of our Institutional Review Board (IRB), we conducted a survey study to investigate the perceptions of professionals of concepts related to trust when using LLM4SE tasks. We chose a survey study over interviews or mixed methods to gather a broader and more diverse range of practitioner insights on trust in LLMs for SE tasks. User studies, particularly surveys, enable efficient data collection from larger samples while maintaining consistency for comparisons between participant groups. Unlike interviews, which are time-intensive and prone to interviewer bias, surveys provide anonymity, reducing social desirability bias, and ensuring honest responses. Furthermore, the inclusion of open-ended questions in the survey allowed us to capture qualitative insights alongside quantitative data, offering a balance between depth and scalability suitable for our exploratory research goals. This approach aligns with established practices in trust research, improving the relevance and generalizability of our findings. In this study, we specifically explore the definitions of trust, its antecedents, and metrics, complementing our systematic review of the literature with insights from experienced and novice practitioners. We employed purposive sampling (Baltes and Ralph, 2021) to recruit practitioners with varying levels of SE and ML expertise, including industrial experts, researchers, students, and software engineers. We reached out to participants using our academic and industrial networks. We also encouraged our network to forward the invitation to relevant colleagues or peers in their networks, thus broadening our reach. Participation was voluntary, no incentives were offered, and participants were informed of the voluntary nature of our study during the request.
@@ 4.2.1. Survey Structure
Section 1 examines the practitioner’s familiarity with using LLMs for various SE tasks. It qualifies practitioners for the rest of the survey, ensuring valid feedback by excluding those who have never used LLMs for subsequent downstream tasks.
Section 2 combines blocks of questions investigating practitioners’ insights into trust-related concepts in Test Case Generation, Code Generation, and Program Repair. The blocks are presented in random order to avoid bias. Each block covers trust definitions, antecedents, and metrics. Only practitioners with experience in using LLMs for these tasks answer these questions.
Section 3 captures the general perception of the importance of trust, the desired metric of trust, and whether practitioners value the importance of these topics.
@@ 4.2.2. Qualitative Metric
  • Definition: We ask practitioners with experience using LLMs for specific tasks to define trust in code generation, test case generation, and program repair.
  • Antecedents of Trust: We gather practitioners’ views on useful trust antecedents, as explained in 4.2.3, for particular downstream SE tasks, which allows us to investigate whether trust antecedents are consistent across downstream tasks.
  • Trust Metrics: We collect information on how practitioners measure trust in specific downstream tasks through open-ended questions. This helps us to compare the trust measurement methods of practitioners with those in the literature. We also ask about their code review process for both human-written code and LLMs generated code to identify similarities and differences.
@@ 4.2.3. Trust Antecedents Selection
The survey asks participants to select trust antecedents for each downstream task. We identified attributes related to the model and the community in the SE, HCI, and ML literature (Li et al., 2023b; EUw, [n. d.]; Körber, 2018; Widder et al., 2021; Johnson et al., 2023; Ji et al., 2024). An author compiled all documented antecedents and then two authors independently categorized them, resolving disagreements through discussion. We selected the nine most popular antecedents for our survey, allowing participants to add any additional ones that were not originally listed.
@@ 4.2.4. Data Collection
We reached out to $96$ potential participants with varying expertise and from different backgrounds, who were not involved or aware of the purpose of this work. Of these groups, $48$ participants answered the survey. However, we had to discard the $23$ participants due to the incomplete or poor quality of their responses leaving $25$ valid responses. The study was performed with Qualtrics (noa, [n. d.]).
@@ 4.2.5. Survey Participants
The survey included 25 participants representing a diverse range of professional roles in software engineering and research. The majority were identified as students (40%), followed by researchers (20%), engineers (16%), and professors (12%), with the remaining roles distributed among freelancers, managers, and other titles. Participants demonstrated various levels of experience in software engineering: 44% reported more than 5 years of experience, 24% had 3 to 5 years and 28% had 1 to 2 years, with only one respondent (4%) having less than a year of experience. In terms of education, the majority had graduate or professional degrees (60%), while 36% had bachelor’s degrees and 4% had completed high school or equivalent. When asked about their experience using LLM for software engineering tasks, 36% of the respondents reported 1 to 2 years of experience, another 36% had less than 1 year, and 16% had 3 to 5 years of experience. This distribution reflects a relatively early stage of adoption of LLMs within software engineering, aligning with the novelty of the technology in this domain. Please refer to our online Appendix (AnonymousRepoTrustRepo, 2024) for more details.
@@ 4.2.6. Survey Validity
To solidify our survey, we conducted two pilot studies: one with $5$ participants and the other with $2$ participants who were not invited to the main survey. Based on the results of these pilot studies, we eliminated any leading questions and minor errors and improved our selection of antecedents of trust. Our survey study is also based on the findings of our literature survey. In addition, two of the authors individually performed the data analysis to avoid bias and human error.
@@ 4.3. Complementary Analysis of Broader Trust literature
Table 1. Representative Papers from Complementary Trust Literature Analysis
Paper Key Contribution Relevance to LLM4SE
Jakesch et al. 2019 (Jakesch et al., 2019) Perception of AI authorship significantly impacts perceived trustworthiness, often leading to distrust in mixed human-AI environments. Highlights need to manage perceptions of LLM involvement to prevent unwarranted distrust of correct LLM-generated code.
Kaur et al. 2022 (Kaur et al., 2022) Comprehensive review of trustworthy AI attributes (robustness, fairness, security) as intrinsic system properties. Provides foundational understanding for conceptualizing trustworthiness objectively in LLM-assisted SE.
Lewis et al. 2018 (Lewis et al., 2018) Synthesizes factors affecting trust and measurement methods, highlighting dynamic trust and calibration issues. Provides foundational understanding of trust dynamics and measurement challenges for managing trust in LLM-assisted SE.
Li et al. 2021 (Li et al., 2023b) Framework for building trustworthy AI across the lifecycle, covering robustness, explainability, accountability. Provides systematic approach for designing, deploying, and governing reliable LLM-based SE tools.
Kim et al. 2024 (Kim et al., 2024) Shows first-person uncertainty expressions reduce overreliance on incorrect outputs and calibrate trust. Highlights importance of explicit communication of LLM limitations to calibrate user trust in SE.
Langer et al. 2021 (Langer et al., 2021) Conceptual model linking explainability to satisfaction of diverse stakeholder needs via human understanding. Emphasizes considering stakeholder-specific desiderata beyond technical performance when building trust in LLMs.
Huang et al. 2024 (Huang et al., 2024) Principles for eight LLM trustworthiness dimensions and empirical benchmark to evaluate mainstream LLMs. Provides systematic framework for conceptualizing and measuring LLM trustworthiness relevant to SE contexts.
Lee & See 2004 (Lee and See, 2004) Introduces “appropriate reliance” and trust calibration aligning user trust with actual system performance. Foundational model for understanding and managing trust levels in human-LLM interactions.
Vereschak et al. 2021 (Vereschak et al., 2021) Survey of empirical measurement methods highlighting the need for explicit trust definitions in experiments. Guides empirical studies on trust in LLM-assisted SE, emphasizing clear definitions for valid results.
Ye et al. 2023 (Ye et al., 2023) TRUSTLLM framework assessing eight trustworthiness dimensions with empirical evaluation. Provides actionable methods for assessing LLM trustworthiness, improving reliability, safety, and ethics in SE tools.
Our initial review of trust in LLMs in SE yielded only $18$ articles, revealing significant gaps in existing literature in LLM4SE. Thus, to supplement our systematic review and provide additional context for trust in LLM4SE, we performed a complementary analysis of broader trust literature from HCI, DL and Automation. These fields provide well-established definitions, influencing factors, and evaluation metrics that can inform our understanding of trust in LLM4SE. Thus, we extracted trust-related concepts, such as definitions, influencing factors, and evaluation metrics, that could inform our findings within the SE domain. Table 1 highlights a selection of representative papers that exemplify these concepts and their contributions.